# Beyond Direct Generation: A Decomposed Approach to Well-Crafted Screenwriting with LLMs

### v2

@@ 1 Introduction
Current research has made notable progress in controllable story generation, developing models that follow plot outlines (rashkin2020plotmachinesoutlineconditionedgenerationdynamic) or incorporate commonsense knowledge (wang2022incorporatingcommonsenseknowledgestory). In screenwriting, systems like Dramatron (mirowski2022cowritingscreenplaystheatrescripts) employ hierarchical prompt chaining.
@@ 2 Related Works
Screenplay Generation. Early screenplay generation relied on retrieval-based methods (10.1145/3507356). Modern LLM-based frameworks like Dramatron (mirowski2022cowritingscreenplaystheatrescripts) and IBSEN (han2024ibsendirectoractoragentcollaboration) enable interactive co-writing but require substantial human supervision. tian2024largelanguagemodelscapable found that while LLMs excel at surface-level qualities, they fail at "narrative intelligence" in end-to-end generation due to the difficulty of simultaneously managing creative construction and structural constraints. Unlike collaborative frameworks, our work tackles autonomous, high-quality generation.
Planning-based Narrative Generation. Plan-and-write approaches use sparse plans (brei2023returning; wang2024guiding; wang2023improving) or detailed outlines (yang2023docimprovinglongstory), but these create semantic gaps between planning and execution. Sophisticated pipelines like CML-BENCH (zheng2025cmlbenchframeworkevaluatingenhancing), HoLLMwood (kor2023hollmwood), and R2 (lin2025r) still rely on lossy, non-narrative representations. Iterative refinement methods like Re3 (yang2022re3generatinglongerstories) attempt costly post-generation fixes.
@@
Data Synthesis for Narrative Tasks. High-quality training data is scarce for screenplay generation. Prior work employs reverse synthesis (you2023eipe; huang2023ex3; ahuja2024finding) to extract plans from texts, or forward synthesis (yang2024rlcd; zhu2024end) using LLMs to generate training data. However, these focus on single-stage generation.
@@ 3.2 The DSR Framework
The choice of novelistic prose as the intermediate representation is a central component of our framework’s design. This choice is motivated by two observations about LLMs (openai2024gpt4technicalreport; touvron2023llamaopenefficientfoundation). First, while LLMs are pre-trained on vast corpora of narrative texts such as novels and stories, they have seen comparatively limited well-formatted screenplay data during training. Second, there exists a significant difference in information density between the two formats: novels can unfold plots with extensive descriptions and details, whereas screenplays must compress the same narrative into a much more concise format with strict structural constraints. Therefore, we introduce the novel as an intermediate representation, decomposing the generation task into two subtasks: first generating a novel from the outline, then converting the novel into a screenplay.

### v3

@@ 1 Introduction
Current research has made notable progress in controllable story generation, developing models that follow plot outlines (Rashkin et al., 2020) or incorporate commonsense knowledge (Wang et al., 2022). In screenwriting, systems like Dramatron (Mirowski et al., 2022) employ hierarchical prompt chaining.
@@ 2 Related Works
Screenplay Generation. Early screenplay generation relied on retrieval-based methods (Zhu et al., 2022). Modern LLM-based frameworks like Dramatron (Mirowski et al., 2022) and IBSEN (Han et al., 2024) enable interactive co-writing but require substantial human supervision. Tian et al. (2024) found that while LLMs excel at surface-level qualities, they fail at "narrative intelligence" in end-to-end generation due to the difficulty of simultaneously managing creative construction and structural constraints. Unlike collaborative frameworks, our work tackles autonomous, high-quality generation.
Planning-based Narrative Generation. Plan-and-write approaches use sparse plans (Brei et al., 2023; Wang and Kreminski, 2024; Wang et al., 2023) or detailed outlines (Yang et al., 2023b), but these create semantic gaps between planning and execution. Sophisticated pipelines like CML-BENCH (Zheng et al., 2025), HoLLMwood (Kor et al., 2023), and R2 (Lin et al., 2024) still rely on lossy, non-narrative representations. Iterative refinement methods like Re3 (Yang et al., 2022) attempt costly post-generation fixes.
@@
Data Synthesis for Narrative Tasks. High-quality training data is scarce for screenplay generation. Prior work employs reverse synthesis (You et al., 2023; Huang et al., 2023; Ahuja et al., 2024) to extract plans from texts, or forward synthesis (Yang et al., 2023a; Zhu et al., 2024) using LLMs to generate training data. However, these focus on single-stage generation.
@@ 3.2 The DSR Framework
The choice of novelistic prose as the intermediate representation is a central component of our framework’s design. This choice is motivated by two observations about LLMs (OpenAI et al., 2024; Touvron et al., 2023). First, while LLMs are pre-trained on vast corpora of narrative texts such as novels and stories, they have seen comparatively limited well-formatted screenplay data during training. Second, there exists a significant difference in information density between the two formats: novels can unfold plots with extensive descriptions and details, whereas screenplays must compress the same narrative into a much more concise format with strict structural constraints. Therefore, we introduce the novel as an intermediate representation, decomposing the generation task into two subtasks: first generating a novel from the outline, then converting the novel into a screenplay.